[ the hypotenuse of a triangle measures 13 mm and one of the other sides measures 5mm. determine the area of the triangle. ]
Answer by TimothyLamb(4379)

h = sqrt( aa + bb )
---
h = sqrt( aa + bb )
13 = sqrt( aa + 5*5 )
13^2 = aa + 25
aa = 144
---
a = 12
b = 5
---
area = (1/2)base*height
---
Answer:
area = (1/2)12*5
area = 30 sq.mm
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php